OpenStudy (anonymous):

lim [(x^2+3x-1)tanx]/(x^2+2x) . x->0

OpenStudy (anonymous):

= (1 + 3/x - 1/x^2)tanx/(1 + 2/x) = (1 + 0 - 0)*0/(1 + 0) = 0

OpenStudy (anonymous):

Answer is supposed to be -0,5

OpenStudy (zarkon):

note that \[\lim_{x\to 0}\frac{\tan(x)}{x}=1\]

OpenStudy (alfie):

see the attachment ^

OpenStudy (alfie):

First step I multiplied and divided everything by "x" so that I get the limit which Zarkon here said ;)

OpenStudy (anonymous):

Thanks!

OpenStudy (alfie):

You're welcome :)

OpenStudy (anonymous):

damn. where did i go wrong? i can't see it

OpenStudy (alfie):

You have a division by zero there, check the 1+2/x

OpenStudy (anonymous):

oh wait, i'm a complete idiot. i evaluated the limit of 2/x to be 0...